Anuradha Sharma, a start up entrepreneur from Bareilly, has invested Rs 80 lacs in an apparel retail



store. Business has been good, and the store shows an accounting profit of Rs 10 lacs for the last year.



This profit is after taxes and after payment of a Rs 20 lacs salary to Ms. Sharma. This salary is less than



what she could make at another job, which is about equal to Rs 40 lacs. Considering the risk involved in



the fashion retail business post Covid’19, she believes that a 15 percent after-tax rate of return is



appropriate for this type of investment. (20 marks)



a. Given this information, calculate the economic profit earned by Ms Sharma.



b. What accounting profit would the firm have to earn in order for the firm to break even in term



of economic profit?




1
Expert's answer
2021-12-06T16:40:45-0500

a.

"EP=IC(ROIC-WACC)=80(0.15-\frac{10}{80})=2"

b.


 "R=\frac{Profit}{Expenses}"


"0.15=\frac{10}{Expenses}"


Expenses=66.67


Expenses=66.67+20=86.67 (40-20=20)

R=0.15


 "0.15=\frac{Profit}{86.67}"


Profit=13